### v1

# Deep Priors for Video Quality Prediction

Abstract.
In this work, we designed a completely blind video quality assessment algorithm using the deep video prior. This work mainly explores the utility of deep video prior in estimating the visual quality of the video. In our work, we have used a single distorted video and a reference video pair to learn the deep video prior. At inference time, the learned deep prior is used to restore the original videos from the distorted videos. The ability of learned deep video prior to restore the original video from the distorted video is measured to quantify distortion in the video. Our hypothesis is that the learned deep video prior fails in restoring the highly distorted videos. The restoring ability of deep video prior is proportional to the distortion present in the video. Therefore, we propose to use the distance between the distorted video and the restored video as the perceptual quality of the video. Our algorithm is trained using a single video pair and it does not need any labelled data. We show that our proposed algorithm outperforms the existing unsupervised video quality assessment algorithms in terms of LCC and SROCC on a synthetically distorted video quality assessment dataset.
copyright: rightsretaineddoi: 10.475/123_4isbn: 123-4567-24-567/08/06conference: 14th Indian Conference on Computer Vision, Graphics and Image Processing; December 2024; Bangalore, Indiaccs: Computing methodologies Image processingccs: Computing methodologies Unsupervised learningccs: General and reference Evaluationccs: General and reference Performance
1. Introduction
Video content and the scale of video consumption has been increasing due to high speed internet and mobile devices. Good objective video quality assessment (VQA) algorithms are required to efficiently manage this content. Objective VQA algorithms are classified into three categories: i) Full-Reference (FR) ii) Reduced-Reference (RR) and iii) No-Reference (NR) methods. FR-VQA algorithms require distorted video and also the corresponding reference video. RR-VQA algorithms require partial knowledge of the reference video while NR-VQA algorithms do not need any knowledge of the reference video for predicting the quality of the test video. We can classify the NR-VQA algorithms into sub categories: i) Supervised methods, which use subjective quality scores as labels while designing the algorithm. ii) Unsupervised methods (or blind methods), which do not use any subjective quality scores (labelled data) while designing the algorithm.
The main drawback of the Supervised NR-VQA algorithms is generalizability across the distortion types and also across the datasets. So it is important to design a universal video quality predictor, which is not biased towards datasets. There are few unsupervised NR-VQA algorithms (Mittal et al., 2012, 2015; Tu et al., 2020) in the literature. The unsupervised video quality algorithms VIIDEO (Mittal et al., 2015) and NIQE (Mittal et al., 2012) build on the statistical priors learned from natural videos and natural images respectively. The low level statistics of images and videos are modelled using probabilistic models. The Natural Image Quality Evaluator (NIQE) (Mittal et al., 2012) assumes that all the natural images posses a unique statistical signature. The mean subtracted contrast normalised coefficients (MSCN) of the natural images follows a Gaussian distribution (Ruderman and Bialek, 1994). However, the distribution of MSCN coefficients of the distorted images do not follow a Gaussian distribution. The statistical prior model describing the set of natural images is estimated. In order to get the visual quality of the distorted image, NIQE (Mittal et al., 2012) computes the distance between the statistical model of the distorted image and the statistical prior estimated from the corpus of natural images. In order to get the video level quality, mean of NIQE scores of all the frames is computed. The completely blind video oracle (Mittal et al., 2015) models the MSCN coefficients of frame differences using a statistical model. The statistical model estimated from the set of natural videos is used as a prior. To estimate a test video’s quality, it measures how far is the test video’s statistical model from the reference statistical model.
Statistical priors have been used thus far in designing the video quality assessment algorithms. Recently, frameworks such as Deep Image Prior (DIP) (Ulyanov et al., 2018) and Deep Decoder (DD) (Heckel and Hand, 2018) have been proposed to capture image priors. These frameworks claim that the structure of a convolutional neural network (CNN) is sufficient to capture the low level image statistics. The CNN structure itself acts as a prior. The Deep Video Prior (DVP) model (Lei et al., 2020) is proposed to impose temporal consistency while learning the model on a pair of original and processed frames. In general, natural videos are smooth and temporally consistent. In the literature, temporal consistency is enforced explicitly computing optical flow or motion vectors while solving video tasks. The DVP model (Lei et al., 2020) claims that the feature maps of a CNN on two similar patches are expected to be similar at the early stage of training. The objective function of DVP is proposed to ensure that CNN feature maps of the predicted frames and the ground truth frames are similar. This kind of objective function is sufficient to impose temporal consistency. This captures the smoothness present in natural videos. So, the model learned using this objective function acts as the prior describing the natural videos. All these deep prior models (Ulyanov et al., 2018),(Heckel and Hand, 2018) and (Lei et al., 2020) show that excellent performance can be achieved by using only test data. Unlike most previous methods that enforce temporal consistency with optical flow, this method shows that the temporal consistency can be achieved by training a convolutional network on a video with the Deep Video Prior. We have used this deep prior model to design an blind video quality algorithm. These models are performing well in many computer vision tasks like super resolution and denoising and enhancement.
2. Proposed Method
Figure 1. Training the Deep Video Prior
Given the efficacy of the DVP (Lei et al., 2020) in video colorization and video enhancement tasks, we propose to use this model in designing an unsupervised video quality assessment algorithm. A Fully Convolutional Network $(G)$ is used to learn the distortion operator as shown in fig. 1. The main objective of the model $G$ is to restore the original frames from the given distorted frames. The temporal consistency is enforced through loss function. $\{D_{t}\}_{t=1}^{T}$ represents the distorted video and $\{O_{t}\}_{t=1}^{T}$ represents the original video. The distorted frames $\{D_{t}\}_{t=1}^{T}$ are passed through the model $G$ to get the restored frames $\{R_{t}=G(D_{t})\}_{t=1}^{T}$. We train the model such that the restored frames $\{R_{t}\}_{t=1}^{T}$ look like the original frames $\{O_{t}\}_{t=1}^{T}$. We have used the objective function described in equation 1 to update the weights $\theta$ of the network $G$. The loss function $L_{\text{perceptual}}$ (1) computes the $L_{1}$ norm between the features of restored frame $R_{t}$ and the original frame $O_{t}$ at each layer of the network $G$. $F_{k}$ indicates the CNN features of the $k^{th}$ layer. We have used a Fully Convolutional Network based on VGG-19 architecture for model $G$. A single video pair (the distorted video $\{D_{t}\}_{t=1}^{T}$ and the original video $\{O_{t}\}_{t=1}^{T}$ is used to train the model $G$. The model $G$ is initialized randomly and trained for 10 epochs.
(1) $L_{\text{perceptual}}(G(D_{t}),O_{t})=\min_{\theta\epsilon\mathcal{G}}\sum_{k=1}^{L}||F_{k}(G(D_{t})-F_{k}(O_{t})||_{1}$
2.1. Quality Prediction
Generally, the natural videos are inherently smooth and temporally consistent. Therefore, quantifying the temporal inconsistency gives us the perceptual quality score of a video The learned model $G$ is further used to predict the quality on the test videos. We pass the test videos (distorted videos) through the model $G$. We predict the visual quality of the distorted video by measuring how well the deep video prior is restoring the original frames from the distorted frames. We build on the fact that highly distorted videos are not restored well using the deep video prior. The distance between distorted frames and restored frames is used as quality of the video. The higher distortion in the video leads to artifacts in restored frames as well. So the model $G$ fails in reconstructing the original frames well. We claim that the distance between restored frames and distorted frames is proportional to visual quality of the video. The equation 2 describes the computation of the video quality score.
(2) Quality $\text{Quality Score}=\frac{1}{T}\sum_{t=1}^{T}\text{log}({\text{PSNR}}(R_{t},D_{t})).$
The mean of the log of peak signal to noise ratio (PSNR) between restored frames and distorted frames is used as he quality score of the video. $T$ is the number of frames in the given test video.
3. Results and Discussion
We used the ECVQ dataset (Vranješ et al., 2012) to show the improved performance of our quality prediction algorithm. It has 90 videos of resolution $352\times 288$. The Linear Correlation Coefficient (LCC) and Spearman Rank Order Correlation Coefficient (SROCC) scores are used to evaluate the video quality assessment algorithms. The LCC and SROCC are computed between the subjective scores of the videos and predicted video quality scores. The higher values of LCC and SROCC indicates that the proposed algorithm is correlating well with the human perpetual quality. We compared our proposed method with the completely blind algorithms. Our results are reported in the table 1. Our proposed method is outperforming both the methods, which build on the statistical priors.
Method LCC SROCC
NIQE (Mittal et al., 2012) 0.4960 0.4469
VIIDEO (Mittal et al., 2015) 0.28 0.15
Proposed Method 0.5089 0.5209
Table 1. Results on ECVQ Dataset (Hosu et al., 2017) for completely blind video quality assessment
4. Conclusion
Through this preliminary attempt, we demonstrate that DVP can be used in designing a video quality predictor. To the best of our knowledge, this is the first work that successfully applies the deep learning based prior to the video quality prediction. Our method uses only a single reference and test video pair for training the DVP. The models (Lei et al., 2020; Heckel and Hand, 2018; Ulyanov et al., 2018) have been trained using single test sample are shown to be performing well in other computer vision tasks like denoising and colorization etc. However, these frameworks have not been applied yet to quality assessment tasks. We believe that this opens up a new direction for video or image quality assessment research to use the deep video and image priors. As future work, we plan to build on these preliminary results by using more advanced image priors on a frame wise basis. We would like to extend this approach further to design an efficient unsupervised stand-alone approach for video quality prediction.